# The Bayesian Origin of the Probability Weighting Function in Human Representation of Probabilities

### v1

@@ 1 Introduction
It is a long-standing observation that human representation of probability is distorted. In decision-making under risk, this manifests as systematic deviations from the Expected Utility framework, as highlighted by the Allais paradox (Allais, 1953). Prospect Theory (Kahneman & Tversky, 1979) addressed these deviations by introducing a probability weighting function, typically inverse S-shaped: small probabilities are overweighted, while large probabilities are underweighted (Figure 1A). This function has been central in explaining a wide range of behavioral anomalies (Ruggeri et al., 2020).
However, a fundamental question remains unanswered: what is the origin of the probability weighting function? Classical approaches have proposed parametric forms (e.g. Prelec, 1998; Zhang & Maloney, 2012), which describe but do not explain its shape. More recent work (e.g. Fennell & Baddeley, 2012; Steiner & Stewart, 2016; Zhang et al., 2020; Khaw et al., 2021; Frydman & Jin, 2023; Bedi et al., 2025; Enke & Graeber, 2023) suggests that, rather than following a deterministic transformation, probabilities are imprecisely encoded in the mind, and that properties of this encoding give rise to distortions, for instance because humans combine noisy encodings with prior expectations about the value of probabilities. However, the specific process by which the S-shaped distortion results remains unclear, with ideas such as log-odds-based transformations (Zhang et al., 2020; Khaw et al., 2021), biases away from the bounds of the response range (Fennell & Baddeley, 2012; Bedi et al., 2025), or efficient coding (Frydman & Jin, 2023).
In this paper, we show that probability representation can be parsimoniously explained in terms of optimal decoding from noisy internal representations in the brain (Figure 1B). In our Bayesian framework, probabilities are imprecisely encoded and decoded via Bayes risk minimization, naturally giving rise to systematic distortions in perceived probability. This account not only unifies prior proposals as special cases but also yields testable theoretical predictions. Most importantly, we demonstrate analytically that the widely observed inverse S-shaped weighting function implies a U-shaped allocation of encoding resources. We then test this and related predictions across domains—perceptual judgment, economic decision-making under risk, and adaptation to novel stimulus statistics—and show that the Bayesian framework consistently provides closer quantitative fits in head-to-head comparison with competing models. Overall, our results show that probability distortion arises from Bayesian decoding of noisy representations, rather than from arbitrary transformation of probabilities.
@@ 2 Background and Relevant work
A large body of work in economics and psychology models probability distortion with parametric forms of the probability weighting function. Classical examples include the inverse-S shaped forms in Prospect Theory (Tversky & Kahneman, 1992), the Prelec function (Prelec, 1998), and the Linear-in-Log-Odds (LILO) model (Zhang & Maloney, 2012), which assumes that under the log-odds transform, the perceived probability $\widehat{p}$ is linear in the true probability $p$ ($\gamma,\beta$ are free parameters):
@@
More recent work has shifted focus from deterministic transformations to the idea that probabilities are encoded imprecisely in the brain. Under this view, distortions arise as systematic consequences of noisy internal representations. Several mechanisms have been proposed. Regression-based accounts emphasize biases away from the boundaries of the probability scale (Fennell & Baddeley, 2012; Bedi et al., 2025). Log-odds based approaches assume that internal representations of probability are approximately linear in log-odds space (Khaw et al., 2021; 2022). Other recent work highlights efficient-coding accounts (Frydman & Jin, 2023) and optimal inference under noisy encodings (Juechems et al., 2021; Enke & Graeber, 2023). Together, these models provide potential explanatory foundations for probability weighting, but they diverge in their specific assumptions.
The Bounded Log-Odds (BLO) model (Zhang et al., 2020) has emerged as a leading explanatory account within the noisy-encoding class. BLO assumes that probability $p$ is first mapped to log-odds and truncated to a bounded interval $\Delta_{-},\Delta_{+}]$2). The clipped value is then linearly mapped to an encoding $\Lambda(p)$ on an interval $\Psi,\Psi]$3), and then combined with an “anchor” $\Lambda_{0}$, where $\kappa>0$ is a free parameter( 4). This encoding is subject to Gaussian noise and then decoded into an estimate of the probability by applying the inverse log-odds function (5).
$\lambda(p)$ $=\log\tfrac{p}{1-p},\qquad\Gamma(\lambda)=\min{(\max(\lambda,\Delta_{-}),\Delta_{+}})$ (2)
@@
In sum, parametric models describe the distortion’s shape but lack explanation, while noisy-encoding models offer explanations but with divergent assumptions. Our Bayesian framework unifies these perspectives within a normative account, deriving general predictions (Section 3.1) that we test empirically.
@@ 3 A general Bayesian framework for perceived probability
The traditional approach assumes that probabilities are deterministically transformed, using a domain-specific, nonlinear function (Tversky & Kahneman, 1992). Here, we propose a different account based on general principles of perceptual processing. Specifically, we model the probability $p$ as encoded into a noisy internal signal $m$, from which the decision maker derives an estimate $\hat{p}$ by minimizing Bayes risk. Distortions of probabilities arise because the optimal Bayesian estimate is typically biased (e.g. Knill & Richards, 1996; Weiss et al., 2002; Körding & Wolpert, 2004). This encoding-decoding approach has been successful in accounting for biases in the perception of numerosity, orientation, color, time intervals, and subjective value (e.g. Polanía et al., 2019; Summerfield & De Lange, 2014; Fritsche et al., 2020; Woodford, 2020; Jazayeri & Shadlen, 2010; Girshick et al., 2011; Wei & Stocker, 2015; 2017; Hahn & Wei, 2024).
@@
where $F:[0,1]\rightarrow\overline{\mathbb{R}}$ is a general, strictly monotone increasing and smooth function. $m$ represents the neural population code for $p$ in the brain. The slope of $F(p)$ is the encoding precision, with Fisher Information (FI) $\mathcal{J}(p)=(F^{\prime}(p))^{2}/\sigma^{2}$, i.e., greater slope of $F$ indicates greater encoding precision. Following Hahn & Wei (2024), we refer to $\sqrt{\mathcal{J}(p)}$ as the (encoding) resources allocated to $p$. Given $m$ and a prior distribution $P_{prior}(p)$, Bayesian inference yields a posterior $P(p|m)$. We assume that a point estimate $\hat{p}(m)$ is derived as the posterior mean, which minimizes the Bayes risk for the mean-squared loss function. The model’s behavior is therefore governed by the encoding $F$ and the prior $P_{\text{prior}}$, both of which we will infer from behavioral data.
Multiple existing accounts represent special cases of this framework (Fennell & Baddeley, 2012; Frydman & Jin, 2022; Bedi et al., 2025; Khaw et al., 2021; 2022; Enke & Graeber, 2023), such as with a uniform FI (Fennell & Baddeley, 2012), a log-odds-based encoding (Khaw et al., 2021; 2022), or efficient coding (Frydman & Jin, 2022). We will discuss these accounts in Section 3.1.2.
@@ 3.1.1 The Predictions of Our Framework
We first derive the estimation bias under the Bayesian framework (see Appendix B.1.1 for the proof):
@@
At any $p\in(0,1)$, the Bayesian model has the bias, assuming $F(0),F(1)$ are finite:
$\underbrace{\operatorname{sign}(0.5-p)\frac{A_{1,\sigma}(p)}{\sqrt{\mathcal{J}(p)}}}_{\text{Regression from Boundary}}+\underbrace{(1-A_{2,\sigma}(p))\left(\frac{1}{\mathcal{J}(p)}\right)^{\prime}}_{\text{Likelihood Repulsion}}+\underbrace{(1-A_{3,\sigma}(p))\frac{1}{\mathcal{J}(p)}\left(\log P_{{prior}}(p)\right)^{\prime}}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$ (8)
as $\sigma\rightarrow 0$. where $A_{\dots,\sigma}(p)$ are bounded, positive, and converge to 0 as $\min\{|F(p)-F(0)|/\sigma,|F(p)-F(1)|/\sigma\}$ increases. Second, if $F(0)=$-$\infty$, $F(1)=\infty$, the theorem remains valid with $A_{\dots,\sigma}(p)\equiv 0$.
The bias arises from three sources, using the terminology of Hahn & Wei (2024): regression away from the boundaries 0 and 1 when the endpoint precision is finite (first term); a repulsion away from high-FI stimuli (second term), and attraction towards regions of high prior density (third term). We first consider the first two terms, which are determined by the Fisher information. The probability weighting function $w(p)$ is typically taken to be S-shaped and to have fixed points at the boundaries, i.e., $w(0)=0$ and $w(1)=1$ (Figure 1A) (Prelec, 1998; Tversky & Kahneman, 1992).
Figure 2: Impact of encoding on bias. (A) U-shaped (blue) and uniform (red) encoding resources $\sqrt{\mathcal{J}(p)}$. (B) The U-shaped encoding generates an S-shaped Likelihood Repulsion bias, $\propto(1/\mathcal{J}(p))^{\prime}$. (C) The uniform encoding generates a Regression bias $\propto\operatorname{sign}(0.5-p)\frac{A_{1,\sigma}(p)}{\sqrt{\mathcal{J}(p)}}$, maximized at 0,1. Both biases generate overestimation of small and underestimation of large probabilities, but with distinct shapes.
For the bias to vanish at the boundaries, the Regression term (Figure 2C) must vanish, which requires $\mathcal{J}(0),\mathcal{J}(1)$ approaching infinity. As infinite precision is biologically implausible in neural populations, we predict a very high but finite FI, leading to near-zero bias at the endpoints. If Regression is neutralized, the Likelihood Repulsion term $\propto(1/\mathcal{J}(p))^{\prime}$ remains; it is S-shaped (positive for small $p$, negative for large $p$, Figure 2B) if and only if $\mathcal{J}(p)$ is U-shaped (Figure 2A). We overall thus predict U-shaped Resources:
@@ 3.1.2 Divergent Predictions from Alternative Models
We next contrast Predictions 1–4 with the implications of alternative models. Prediction 1 is shared with models that assume a log-odds encoding (Zhang et al., 2020; Khaw et al., 2021; 2022), other encodings with U-shaped FI (Enke & Graeber, 2023), or efficient coding for a U-shaped prior (Frydman & Jin, 2023). In this case, the nonuniformity of the encoding makes the Repulsion term (Figure 2B) central. In contrast, Fennell & Baddeley (2012) and Bedi et al. (2025) attribute the S-shaped probability weighting entirely to the boundary-induced regression (Figure 2C), potentially modulated by the Prior Attraction. The Regression effect is strongest at 0 and 1 (unless $\mathcal{J}(p)$ approaches infinity there, in which case it vanishes), i.e., such accounts predict the distortion to be particularly large at the boundaries (Figure 2C), which conflicts the S-shape traditionally assumed. Frydman & Jin (2023) propose that the encoding is optimized for mutual information under a prior, which leads to $\sqrt{\mathcal{J}(p)}\propto p_{prior}(p)$ (see Appendix B.2). They note that, if the prior is U-shaped, the resulting model would produce an S-shaped bias. However, they do not explicitly recover the encoding from the behavioral data; we’ll be able to do this on several datasets, and find that prior and encoding are not always matched in behavioral data.
Prediction 2 is shared with models incorporating Bayesian decoding (Fennell & Baddeley, 2012; Bedi et al., 2025); the BLO model allows attraction to the anchor $\Lambda_{0}$ but no general priors. In contrast, if prior and encoding are matched by efficient coding as $\sqrt{\mathcal{J}(p)}\propto p_{prior}(p)$, the posterior mean is biased away from the prior mode (Wei & Stocker, 2015).
Prediction 3 is a general prediction of Bayesian decoding; in contrast, in BLO, response variance is mainly determined by $\Delta_{-},\Delta_{+}$ and $\sigma$, but not the anchor $\Phi_{0}$ – diverging from the Bayesian model (Appendix, Theorem 10).
Prediction 4 about the optimality is again shared by models assuming Bayesian decoding; BLO differs here as well: even when $\sigma,\kappa=0$, the MSE of its decoded estimate is nonzero everywhere in $(0,1)$, entailing suboptimal estimation.
@@ 3.2 Bayesian Models with Log-Odds Encoding
The Bayesian model can accommodate a general mapping $F$ and we will be able to infer it from data. Here, we justify the popular choice of a log-odds mapping (as assumed by Khaw et al. (2021; 2022); Zhang et al. (2020)), and how this leads to a reinterpretation of BLO as an approximation of the Bayesian model. Suppose that observers encode positive counts ($\kappa_{+}$) and negative counts ($\kappa_{-}$):
@@
where $p=\frac{\kappa_{+}}{\kappa_{+}+\kappa_{-}}$. Based on research on magnitude perception, we take $F_{num}$ to be consistent with Weber’s law by assuming the form (e.g. Petzschner & Glasauer, 2011; Dehaene, 2003):
@@
where $\alpha>0$ prevents an infinite FI at zero (Petzschner & Glasauer, 2011), and $N=\kappa_{+}+\kappa_{-}$. What is an optimal 1D encoding $m_{1D}\in\mathbb{R}$ of the rate $p=\frac{\kappa_{+}}{\kappa_{+}+\kappa_{-}}$? We focus on linear encodings with coefficients $w_{1},w_{2}$:

### v2

@@ 1 Introduction
It is a long-standing observation that human representation of probability is distorted. In decision-making under risk, this manifests as systematic deviations from the Expected Utility framework, as highlighted by the Allais paradox (allais1953comportement). Prospect Theory (kai1979prospect) addressed these deviations by introducing a probability weighting function, typically inverse S-shaped: small probabilities are overweighted, while large probabilities are underweighted (Figure 1A). This function has been central in explaining a wide range of behavioral anomalies (ruggeri2020replicating).
However, a fundamental question remains unanswered: what is the origin of the probability weighting function? Classical approaches have proposed parametric forms (e.g. prelec1998probability; ZM2012), which describe but do not explain its shape. More recent work (e.g. fennell2012uncertainty; zhang2020; khaw2021cognitive; frydman2023source; bedi2025probability; enke2023cognitive) suggests that, rather than following a deterministic transformation, probabilities are imprecisely encoded in the mind, and that properties of this encoding give rise to distortions, for instance because humans combine noisy encodings with prior expectations about the value of probabilities. However, the specific process by which the S-shaped distortion results remains unclear, with ideas such as log-odds-based transformations (zhang2020; khaw2021cognitive), biases away from the bounds of the response range (fennell2012uncertainty; bedi2025probability), or efficient coding (frydman2023source).
In this paper, we show that probability representation can be parsimoniously explained in terms of optimal decoding from noisy internal representations in the brain (Figure 1B). In our Bayesian framework, probabilities are imprecisely encoded and decoded via Bayes risk minimization, naturally giving rise to systematic distortions in perceived probability. This account not only unifies prior proposals as special cases but also yields testable theoretical predictions. Specifically, our contributions are two-fold:
  1. 1.
    We theoretically disentangle the predictions made by distinct modeling proposals. By deriving analytical decomposition of bias, we separate the contributions of prior attraction, encoding-induced repulsion and boundary regression, and clarify that these modeling proposals rely on different resource allocations (Section 3).
  2. 2.
    We conduct the first broad empirical evaluation comparing these models against data from three empirical paradigms. Unlike prior work that typically focuses on a single task, we compare five model classes across judgment of relative frequency, pricing, and choice tasks, showing the specific limitations of each account (Section 4).
@@ 2 Background and Relevant work
A large body of work in economics and psychology models probability distortion with parametric forms of the probability weighting function. Classical examples include the inverse-S shaped forms in Prospect Theory (tversky1992advances), the Prelec function (prelec1998probability), and the Linear-in-Log-Odds (LILO) model (ZM2012), which assumes that under the log-odds transform, the perceived probability $\widehat{p}$ is linear in the true probability $p$ ($\gamma,\beta$ are free parameters):
@@
More recent work has shifted focus from deterministic transformations to the idea that probabilities are encoded imprecisely in the brain. Under this view, distortions arise as systematic consequences of noisy internal representations. Several mechanisms have been proposed. Regression-based accounts emphasize biases away from the boundaries of the probability scale (fennell2012uncertainty; bedi2025probability). Log-odds based approaches assume that internal representations of probability are approximately linear in log-odds space (khaw2021cognitive; Khaw2022StateDependent). Other recent work highlights efficient-coding accounts (frydman2023source) and optimal inference under noisy encodings (juechems2021optimal; enke2023cognitive). These models have generally not been broadly evaluated against behavioral data, and their differing predictions are not understood either theoretically or empirically. One related model has already been evaluated on several behavioral datasets and shown success in outperforming fixed probability weighting functions, the Bounded Log-Odds (BLO) model (zhang2020). BLO assumes that probability $p$ is first mapped to log-odds and truncated to a bounded interval $\Delta^{-},\Delta^{+}]$2). The clipped value is then linearly mapped to an encoding $\Lambda(p)$ on an interval $\Psi,\Psi]$3), and then combined with an “anchor” $\Lambda_{0}$, where $\kappa>0$ is a free parameter( 4). This encoding is subject to Gaussian noise and then decoded into an estimate of the probability by applying the inverse log-odds function( 5).
$\lambda(p)$ $=\log\tfrac{p}{1-p},\qquad\Gamma(\lambda)=\min{(\max(\lambda,\Delta^{-}),\Delta^{+}})$ (2)
@@ 3 A general Bayesian framework for perceived probability
The traditional approach assumes that probabilities are deterministically transformed, using a domain-specific, nonlinear function (tversky1992advances). Here, we propose a different account based on general principles of perceptual processing. Specifically, we model the probability $p$ as encoded into a noisy internal signal $m$, from which the decision maker derives an estimate $\hat{p}$ by minimizing Bayes risk. Distortions of probabilities arise because the optimal Bayesian estimate is typically biased (e.g. Knill1996perception; Weiss2002motion; Kording2004bayesian). This encoding-decoding approach has been successful in accounting for biases in the perception of numerosity, orientation, color, time intervals, and subjective value (e.g. polania2019efficient; summerfield2014expectation; fritsche2020bayesian; woodford2020modeling; jazayeri2010temporal; girshick2011cardinal; wei2015bayesian; wei2017lawful; hahn2024unifying).
@@
where $F:[0,1]\rightarrow\overline{\mathbb{R}}$ is a general, strictly monotone increasing and smooth function. $m$ represents the neural population code for $p$ in the brain. The slope of $F(p)$ is the encoding precision, with Fisher Information (FI) $\mathcal{J}(p)=(F^{\prime}(p))^{2}/\sigma^{2}$, i.e., greater slope of $F$ indicates greater encoding precision. Following hahn2024unifying, we refer to $\sqrt{\mathcal{J}(p)}$ as the (encoding) resources allocated to $p$. Given $m$ and a prior distribution $P_{prior}(p)$, Bayesian inference yields a posterior $P(p|m)$. We assume that a point estimate $\hat{p}(m)$ is derived as the posterior mean, which minimizes the Bayes risk for the mean-squared loss function. The model’s behavior is therefore governed by the encoding $F$ and the prior $P_{\text{prior}}$, both of which we will infer from behavioral data.
Multiple existing accounts represent special cases of this framework (fennell2012uncertainty; frydman2022efficient; bedi2025probability; khaw2021cognitive; Khaw2022StateDependent; enke2023cognitive), such as with a uniform FI (fennell2012uncertainty), a log-odds-based encoding (khaw2021cognitive; Khaw2022StateDependent), or efficient coding (frydman2022efficient). None of these have been evaluated broadly against behavioral data, and their differing predictions are not understood theoretically or empirically. We will discuss these accounts in Section 3.1.2.
@@ 3.1.1 The Predictions of Our Framework
Whereas prior work on the functional form of the bias of Bayesian models of perception largely focused on unbounded or circular stimulus spaces, or stimuli far away from a boundary (hahn2024unifying; wei2017lawful; prat2021bias; morais2018power; stocker2006noise), the bounded nature of probabilities requires a treatment fully accounting for the boundary (see Appendix B.1.1 for the proof):
@@
There are functions $A_{1,\sigma},A_{2,\sigma},A_{3,\sigma}:[0,1]\rightarrow\mathbb{R}_{+}$ that are bounded, positive, and satisfy
$A_{\dots,\sigma}(p)\leq\phi\left(\frac{\sigma}{\min\{|F(p)-F(0)|,|F(p)-F(1)|\}}\right)$ (8)
for some nondecreasing function $\phi:\mathbb{R}\rightarrow\mathbb{R}$ with $\lim_{x\downarrow 0}\phi(x)=0$, across all $p\in(0,1);\sigma>0$. Then, at any $p\in(0,1)$, the Bayesian model has the bias, as $\sigma\rightarrow 0$:
$\operatorname{Bias}(p)=\underbrace{A_{1,\sigma}(p)\cdot\frac{\operatorname{sign}(0.5-p)}{\sqrt{\mathcal{J}(p)}}}_{\text{Regression from Boundary}}+\underbrace{(1-A_{2,\sigma}(p))\cdot\frac{\operatorname{d}}{\operatorname{d}p}\left(\frac{1}{\mathcal{J}(p)}\right)}_{\text{Likelihood Repulsion}}$
$+\underbrace{(1-A_{3,\sigma}(p))\cdot\frac{1}{\mathcal{J}(p)}\cdot\frac{\operatorname{d}}{\operatorname{d}p}\left(\log P_{{prior}}(p)\right)}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$
The bias arises from three sources, using the terminology of hahn2024unifying. The first term is regression effect towards $p=0.5$ (first term); a repulsion away from high-FI stimuli (second term), and attraction towards regions of high prior density (third term). All three components are modulated by terms $A_{1,\sigma},A_{2,\sigma},A_{3,\sigma}$ that intuitively represent the effect of the boundary: Eq. 8 expresses that these terms vanish when the noise magnitude $\sigma$ is small in relation to the distance of $F(p)$ from both $F(0)$ and $F(1)$: that is, how close $p$ is to one of the two boundaries in the neural encoding space. In particular, if $F(0)=-\infty,F(1)=\infty$, the $A_{\dots,\sigma}$ factors vanish everywhere. On the other hand, when $p$ is close to a boundary, these factors can substantially affect the bias.
We now explain how this theorem drives key predictions. We first consider the first two terms, which are independent of the prior. The probability weighting function $w(p)$ is typically taken to be S-shaped and to have fixed points at the boundaries, i.e., $w(0)=0$ and $w(1)=1$ (Figure 1A) (prelec1998probability; tversky1992advances).
Figure 2: Impact of encoding on bias. (A) A U-shaped (blue) and uniform (red) encoding allocation resources $\sqrt{\mathcal{J}(p)}$. (B) The U-shaped encoding generates an S-shaped Likelihood Repulsion bias, $\propto\frac{\operatorname{d}}{\operatorname{d}p}(1/\mathcal{J}(p))$. (C) The uniform encoding generates a Regression bias $\propto A_{1,\sigma}(p)\cdot\frac{\operatorname{sign}(0.5-p)}{\sqrt{\mathcal{J}(p)}}$, maximized at 0,1. Both biases generate overestimation of small and underestimation of large probabilities, but with distinct shapes.
For the bias to vanish at the boundaries, the Regression term (Figure 2C) must vanish, which requires $\mathcal{J}(0),\mathcal{J}(1)$ approaching infinity. As infinite precision is biologically implausible in neural populations, we predict a very high but finite $\mathcal{J}(0),\mathcal{J}(1)$, leading to near-zero regression bias at the endpoints. If Regression is neutralized, the Likelihood Repulsion term $\propto\frac{\operatorname{d}}{\operatorname{d}p}(1/\mathcal{J}(p))$ remains; it is S-shaped (positive for small $p$, negative for large $p$, Figure 2B) if and only if $\mathcal{J}(p)$ is U-shaped (Figure 2A). We overall thus predict U-shaped Resources:
@@ 3.1.2 Divergent Predictions from Alternative Models
We next contrast Predictions 1–4 with the implications of alternative models.
Prediction 1 is shared with models that assume a log-odds encoding (zhang2020; khaw2021cognitive; Khaw2022StateDependent), other encodings with U-shaped FI (enke2023cognitive), or efficient coding for a U-shaped prior (frydman2023source). In this case, the nonuniformity of the encoding makes the Repulsion term (Figure 2B) central. In contrast, fennell2012uncertainty and bedi2025probability attribute the S-shaped probability weighting entirely to the boundary-induced regression (Figure 2C), potentially modulated by the Prior Attraction. The Regression effect is strongest at 0 and 1 (unless $\mathcal{J}(p)$ approaches infinity there, in which case it vanishes), i.e., such accounts predict the distortion to be particularly large at the boundaries (Figure 2C), which conflicts the S-shape traditionally assumed. frydman2023source propose that the encoding is optimized for mutual information under a prior, which leads to $\sqrt{\mathcal{J}(p)}\propto p_{prior}(p)$ (see Appendix B.2). They note that, if the prior is U-shaped, the resulting model would produce an S-shaped bias. However, they do not explicitly recover the encoding from the behavioral data; we’ll be able to do this on several datasets, and find that prior and encoding are not always matched in behavioral data.
Prediction 2 is shared with models incorporating Bayesian decoding (fennell2012uncertainty; bedi2025probability); the BLO model allows attraction to the anchor $\Lambda_{0}$ but no general priors. In contrast, if prior and encoding are matched by efficient coding as $\sqrt{\mathcal{J}(p)}\propto p_{prior}(p)$, the posterior mean is biased away from the prior mode (wei2015bayesian).
Prediction 3 is a general prediction of Bayesian decoding; in contrast, in BLO, response variance is mainly determined by $\Delta^{-},\Delta^{+}$ and $\sigma$, but not the anchor $\Phi_{0}$ – diverging from the Bayesian model (Appendix, Theorem 10).
Prediction 4 about the optimality is again shared by models assuming Bayesian decoding. BLO differs here as well: even when $\sigma,\kappa=0$, the MSE of its decoded estimate is nonzero everywhere in $(0,1)$, entailing suboptimal estimation.
@@ 3.2 Bayesian Models with Log-Odds Encoding
The Bayesian model can accommodate a general mapping $F$ and we will be able to infer it from data. Here, we justify the popular choice of a log-odds mapping (as assumed by khaw2021cognitive; Khaw2022StateDependent; zhang2020), and how this leads to a reinterpretation of BLO as an approximation of the Bayesian model. Suppose that observers encode positive counts ($\kappa_{+}$) and negative counts ($\kappa_{-}$):
@@
where $p=\frac{\kappa_{+}}{\kappa_{+}+\kappa_{-}}$. Based on research on magnitude perception, we take $F_{num}$ to be consistent with Weber’s law by assuming the form (e.g. petzschner2011iterative; dehaene2003neural):
@@
where $\alpha>0$ prevents an infinite FI at zero (petzschner2011iterative), and $N=\kappa_{+}+\kappa_{-}$. What is an optimal 1D encoding $m_{1D}\in\mathbb{R}$ of the rate $p=\frac{\kappa_{+}}{\kappa_{+}+\kappa_{-}}$? We focus on linear encodings with coefficients $w_{1},w_{2}$: